OpenStudy (anonymous):

Please Help me ;* What is the equation of the line in slope-intercept form that passes through the point (3, –1) and is perpendicular to the line y = x + 1? Possible answers : y = –x – 2 y = x – 2 y = –x + 2 y = x + 2

OpenStudy (anonymous):

Do you know how to find the slope of the perpendicular line?

OpenStudy (anonymous):

nopee

OpenStudy (anonymous):

slope of a perpendicular line is the negative inverse of the slope of the original line

OpenStudy (anonymous):

so if the original slope is 3, slope of perpendicular is -(1/3)

OpenStudy (anonymous):

so here, y = x + 1 is the original line, so the slope of the perp. line is -(1/1) or just -1

OpenStudy (anonymous):

for equations in a form "y=mx+h", m refers to its slope, which means 1 should be the slope of that perpendicular line *just a tip since jake v8 is answering you:S*

OpenStudy (anonymous):

Shouldn't the slope of the perp line be -1 ?

OpenStudy (anonymous):

Feel free to help @mandyau1023 :)

OpenStudy (anonymous):

So if you know the perp line slope is -1, you already know enough to eliminate 2 of the choices. You need to work out the point-slope equation of the perp line, which has slope -1 and passes through the point they gave you, (3,-1)

OpenStudy (anonymous):



OpenStudy (anonymous):

so the answer is y=-x+2 right? (:

OpenStudy (anonymous):

Sorry, stepped away... looks right to me. (3,-1) does fall on y = -x + 2, and the slope of -1 is the negative inverse of y = x+1 slope which was 1, so you are correct.

OpenStudy (anonymous):

Good work :)